# Gyro-based Neural Single Image Deblurring

### v2

@@ 6 Conclusion
Our method has a few limitations. While our method uses a fixed value for the hyperparameter $M$, the complexity of camera shakes may increase along with the exposure time, which means that a longer exposure time may require a larger value for $M$. Our method does not use the accelerometer, which is also equipped with most smartphones nowadays, and which may provide additional valuable information on the camera motion. Our image deblurring module adopts a relatively simple network architecture compared to recent non-gyro-based single image deblurring approaches, and this may limit the deblurring performance of our method. Extending our work to address these limitations would be an interesting future direction.

### v3

@@ 6 Conclusion
Our method has a few limitations. While our method uses a fixed value for the hyperparameter $M$, the complexity of camera shakes may increase along with the exposure time, which means that a longer exposure time may require a larger value for $M$. Our method does not use the accelerometer, which is also equipped with most smartphones, and which may provide additional valuable information on the camera motion. Our image deblurring module adopts a relatively simple network architecture compared to recent non-gyro-based single image deblurring approaches, and this may limit the deblurring performance of our method. Extending our work to address these limitations would be an interesting future direction.
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grants (RS-2023-NR077065, RS-2023-00280400, RS-2023-00211658, RS-2024-00438532) and by Institute of Information & communications Technology Planning & Evaluation (IITP) grant (RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH)) funded by the Korea government (MSIT). This work was also supported by Samsung Research Funding & Incubation Center of Samsung Electronics under Project Number SRFC-IT1801-52.